OpenStudy (anonymous):

Graph each pair of parametric equations. x = 3 sin3t y = 3 cos3t

OpenStudy (anonymous):

help any body, dont understand how to graph or where to start

OpenStudy (anonymous):

do you have a graphing calculator?

OpenStudy (anonymous):

i do not is there one online i can use

terenzreignz (terenzreignz):

Well, you need to take some time to analyse the problem, see if you can find something familiar... in particular, what is x² ? What is y² ?

OpenStudy (anonymous):

ok so how do i plug it in the calculator

OpenStudy (anonymous):

change the "function" to "parametric" and simply type in your equations.

OpenStudy (anonymous):

= sin3t y = cos3t dx/dt = 3cos(3t) and dy/dt = -3sin(3t) dy/dx = (dy/dt)(dt/dx) = -3sin(3t) / 1 / 3cos(3t) = - 9sin(3t)cos(3t) = - 4.5sin(6t) when t = 2π/9. dy/dx = - 4.5sin(12π/9) = - 4.5sin(4π/3)) but x = sin(2π/3) and y = cos(2π/3) when t = 2π/9 so the line is y - cos(2π/3) / {x - sin (2π/3)} = - 4.5sin(4π/3)

OpenStudy (anonymous):

L = 2 ∫ √(18 - 18(cos²(t) - sin²(t))) dt, over [0, π] L = 2 √18 ∫ √(1 - cos²(t) + sin²(t)) dt, over [0, π] L = 2 √18 ∫ √(sin²(t) + sin²(t)) dt, over [0, π] L = 2 √18 ∫ √(2sin²(t)) dt, over [0, π] L = 2 √18 √2 ∫ √(sin²(t)) dt, over [0, π] L = 2 √36 ∫ sin(t) dt, over [0, π] L = 2(6) [-cos(t)] over [0, π] L = 12 (-cos(π) + cos(0)) L = 12 (1 + 1) L = 12(2) so L= (24)

terenzreignz (terenzreignz):

No, @Kuuttboylife let's not complicate life more than it should be :P x = 3sin 3t y = 3cos 3t Which means x² = 9sin² 3t y² = 9cos² 3t So, add them up x² + y² = 9sin² 3t + 9cos² 3t = 9(sin² 3t + cos² 3t) Trig identities remind us that sin² 3t + cos² 3t = 1, so... x² + y² = 9 And I bet, that's MUCH easier to graph, @tomatooooo @TNNG ? ^.^

OpenStudy (anonymous):

IT's 24 lol

OpenStudy (anonymous):

well it is much more clear @Kuuttboylife

OpenStudy (anonymous):

so i graph 9

terenzreignz (terenzreignz):

NO. You graph x² + y² = 9 Which is a....?

OpenStudy (anonymous):

if you looking for the arc length of the curve it's 24

OpenStudy (anonymous):

function?

terenzreignz (terenzreignz):

Nope, @Kuuttboylife the arclength of this curve is not 24, but rather, the circumference of the circle x² + y² = 9, which is \[\large 6\pi\]

OpenStudy (anonymous):

im sorry i graph it has a circle

terenzreignz (terenzreignz):

Yep :)

terenzreignz (terenzreignz):

A circle with radius... what's the radius? :)

OpenStudy (anonymous):

ahh that make some time lol

OpenStudy (anonymous):

but what would i do for x = 7 sin t + sin 7t y = 7 cos t + cos 7t

OpenStudy (anonymous):

nevermind?

OpenStudy (anonymous):

wait yeah how do you do that

terenzreignz (terenzreignz):

Kinda reached my limit there :)

OpenStudy (anonymous):

i feel you there thnx for you help though

terenzreignz (terenzreignz):

No problem :)